Probability

ICSE Class 7 Mathematics · Selina Solutions Concise Mathematics Class 7 ICSE · 35 solved questions

Exercise
Marks
1Exercise 25(A)

A coin is tossed once. Find the probability of:

(i) getting a head

(ii) not getting a head.

Show SolutionHide Solution↓

When a coin is tossed once, the possible outcomes are H and T.

Total number of possible outcomes = 2

(i) Favourable outcomes for getting a head = H

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Getting a head)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{2}\end{aligned}$$

Hence, the probability of getting a head is $\mathbf{\frac{1}{2}}$.

(ii) Favourable outcomes for not getting a head = T

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Not getting a head)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{2}\end{aligned}$$

Hence, the probability of not getting a head is $\mathbf{\frac{1}{2}}$.

2Exercise 25(A)

A coin is tossed 80 times and the head is obtained 38 times. Now, if a coin is tossed once, what will be the probability of getting:

(i) a tail

(ii) a head?

Show SolutionHide Solution↓

Total number of trials = 80

Number of times head is obtained = 38

Number of times tail is obtained = 80 − 38 = 42

(i) Number of favourable outcomes (Getting a tail) = 42

$$\begin{aligned}\text{P(Getting a tail)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{42}{80} \\[1em] = \frac{21}{40}\end{aligned}$$

Hence, the probability of getting a tail is $\mathbf{\frac{21}{40}}$.

(ii) Number of favourable outcomes (Getting a head) = 38

$$\begin{aligned}\text{P(Getting a head)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{38}{80} \\[1em] = \frac{19}{40}\end{aligned}$$

Hence, the probability of getting a head is $\mathbf{\frac{19}{40}}$.

3Exercise 25(A)

A die is thrown 20 times and the outcomes are noted as shown below:

OutcomesNo. of times
12
23
34
44
53
64

Now if the same die is thrown at random, find the probability of getting:

(i) 3

(ii) a number less than 3

(iii) a number greater than 3.

Show SolutionHide Solution↓

Total number of outcomes = 20

(i) The outcome 3 appears 4 times.

Number of favourable outcomes (Getting 3) = 4

$$\begin{aligned}\text{P(Getting 3)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{4}{20} \\[1em] = \frac{1}{5}\end{aligned}$$

Hence, the probability of getting 3 is $\mathbf{\frac{1}{5}}$.

(ii) The numbers less than 3 are 1 and 2.

Number of times 1 or 2 appears = 2 + 3 = 5

Number of favourable outcomes (Getting a number less than 3) = 5

$$\begin{aligned}\text{P(Getting a number less than 3)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{5}{20} \\[1em] = \frac{1}{4}\end{aligned}$$

Hence, the probability of getting a number less than 3 is $\mathbf{\frac{1}{4}}$.

(iii) The numbers greater than 3 are 4, 5 and 6.

Number of times 4, 5 or 6 appears = 4 + 3 + 4 = 11

Number of favourable outcomes (Getting a number greater than 3) = 11

$$\begin{aligned}\text{P(Getting a number greater than 3)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{11}{20}\end{aligned}$$

Hence, the probability of getting a number greater than 3 is $\mathbf{\frac{11}{20}}$.

4Exercise 25(A)

A survey of 50 boys showed that 21 like tea while 29 dislike it. Out of these boys, one boy is chosen at random. What is the probability that the chosen boy

(i) likes tea

(ii) dislikes tea?

Show SolutionHide Solution↓

Total number of boys = 50

Number of boys who like tea = 21

Number of boys who dislike tea = 29

(i) Number of favourable outcomes (Choosing a boy who likes tea) = 21

$$\begin{aligned}\text{P(Choosing a boy who likes tea)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{21}{50}\end{aligned}$$

Hence, the probability that the chosen boy likes tea is $\mathbf{\frac{21}{50}}$.

(ii) Number of favourable outcomes (Choosing a boy who dislikes tea) = 29

$$\begin{aligned}\text{P(Choosing a boy who dislikes tea)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{29}{50}\end{aligned}$$

Hence, the probability that the chosen boy dislikes tea is $\mathbf{\frac{29}{50}}$.

5Exercise 25(A)

In a cricket match, a batsman hits a boundary 12 times out of the 80 balls he plays. Further, if he plays one more ball, what will be the probability that:

(i) he hits a boundary

(ii) he does not hit a boundary?

Show SolutionHide Solution↓

Total number of balls played = 80

Number of times a boundary is hit = 12

Number of times a boundary is not hit = 80 − 12 = 68

(i) Number of favourable outcomes (Hitting a boundary) = 12

$$\begin{aligned}\text{P(Hitting a boundary)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{12}{80} \\[1em] = \frac{3}{20}\end{aligned}$$

Hence, the probability that he hits a boundary is $\mathbf{\frac{3}{20}}$.

(ii) Number of favourable outcomes (Not hitting a boundary) = 68

$$\begin{aligned}\text{P(Not hitting a boundary)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{68}{80} \\[1em] = \frac{17}{20}\end{aligned}$$

Hence, the probability that he does not hit a boundary is $\mathbf{\frac{17}{20}}$.

6Exercise 25(A)

There are 8 marbles in a bag with numbers from 1 to 8 marked on each of them. What is the probability of drawing a marble with number

(i) 3

(ii) 7?

Show SolutionHide Solution↓

The marbles are numbered 1, 2, 3, 4, 5, 6, 7 and 8.

Total number of possible outcomes = 8

(i) Favourable outcomes for drawing the marble numbered 3 = 3

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Drawing the marble numbered 3)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{8}\end{aligned}$$

Hence, the probability of drawing a marble with number 3 is $\mathbf{\frac{1}{8}}$.

(ii) Favourable outcomes for drawing the marble numbered 7 = 7

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Drawing the marble numbered 7)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{8}\end{aligned}$$

Hence, the probability of drawing a marble with number 7 is $\mathbf{\frac{1}{8}}$.

7Exercise 25(A)

Two coins are tossed simultaneously 100 times and the outcomes are as given below:

OutcomesNo. of times
Two heads (H H)21
Exactly one head (H T or T H)55
No head (T T)24

If the same pair of coins is tossed again at random, find the probability of getting:

(i) two heads

(ii) exactly one head

(iii) no head.

Show SolutionHide Solution↓

Total number of trials = 100

(i) Number of times two heads are obtained = 21

$$\begin{aligned}\text{P(Getting two heads)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{21}{100}\end{aligned}$$

Hence, the probability of getting two heads is $\mathbf{\frac{21}{100}}$.

(ii) Number of times exactly one head is obtained = 55

$$\begin{aligned}\text{P(Getting exactly one head)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{55}{100} \\[1em] = \frac{11}{20}\end{aligned}$$

Hence, the probability of getting exactly one head is $\mathbf{\frac{11}{20}}$.

(iii) Number of times no head is obtained = 24

$$\begin{aligned}\text{P(Getting no head)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{24}{100} \\[1em] = \frac{6}{25}\end{aligned}$$

Hence, the probability of getting no head is $\mathbf{\frac{6}{25}}$.

8Exercise 25(A)

A bag contains 4 white and 6 black balls, all of the same shape and same size. A ball is drawn from the bag without looking into the bag. Find the probability that the ball drawn is:

(i) a black ball

(ii) a white ball

(iii) not a black ball

Show SolutionHide Solution↓

Total number of balls = 4 + 6 = 10

(i) Number of favourable outcomes (Drawing a black ball) = 6

$$\begin{aligned}\text{P(Drawing a black ball)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{6}{10} \\[1em] = \frac{3}{5}\end{aligned}$$

Hence, the probability that the ball drawn is black is $\mathbf{\frac{3}{5}}$.

(ii) Number of favourable outcomes (Drawing a white ball) = 4

$$\begin{aligned}\text{P(Drawing a white ball)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{4}{10} \\[1em] = \frac{2}{5}\end{aligned}$$

Hence, the probability that the ball drawn is white is $\mathbf{\frac{2}{5}}$.

(iii) A ball which is not black must be white.

Number of favourable outcomes (Drawing a ball which is not black) = 4

$$\begin{aligned}\text{P(Drawing a ball which is not black)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{4}{10} \\[1em] = \frac{2}{5}\end{aligned}$$

Hence, the probability that the ball drawn is not black is $\mathbf{\frac{2}{5}}$.

9Exercise 25(A)

In a single throw of a die, find the probability of getting a number:

(i) 4

(ii) 6

(iii) greater than 4

Show SolutionHide Solution↓

In a single throw of a die, the possible outcomes are 1, 2, 3, 4, 5 and 6.

Total number of possible outcomes = 6

(i) Favourable outcomes for getting 4 = 4

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Getting 4)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{6}\end{aligned}$$

Hence, the probability of getting 4 is $\mathbf{\frac{1}{6}}$.

(ii) Favourable outcomes for getting 6 = 6

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Getting 6)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{6}\end{aligned}$$

Hence, the probability of getting 6 is $\mathbf{\frac{1}{6}}$.

(iii) The numbers greater than 4 are 5 and 6.

Number of favourable outcomes (Getting a number greater than 4) = 2

$$\begin{aligned}\text{P(Getting a number greater than 4)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{2}{6} \\[1em] = \frac{1}{3}\end{aligned}$$

Hence, the probability of getting a number greater than 4 is $\mathbf{\frac{1}{3}}$.

10Exercise 25(A)

Hundred identical cards are numbered from 1 to 100. The cards are well shuffled and then a card is drawn. Find the probability that the number on the card drawn is:

(i) 50

(ii) 80

(iii) 40

Show SolutionHide Solution↓

Total number of cards = 100

Since the cards are numbered from 1 to 100, there is exactly one card bearing each number.

(i) Favourable outcomes for drawing card number 50 = 50

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Drawing card number 50)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{100}\end{aligned}$$

Hence, the probability of drawing card number 50 is $\mathbf{\frac{1}{100}}$.

(ii) Favourable outcomes for drawing card number 80 = 80

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Drawing card number 80)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{100}\end{aligned}$$

Hence, the probability of drawing card number 80 is $\mathbf{\frac{1}{100}}$.

(iii) Favourable outcomes for drawing card number 40 = 40

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Drawing card number 40)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{100}\end{aligned}$$

Hence, the probability of drawing card number 40 is $\mathbf{\frac{1}{100}}$.

1Exercise 25(B)

Suppose S is the event that it will snow tomorrow and P(S) = 0.03.

(i) State, in words, the complementary event S′.

(ii) Find P(S′)

Show SolutionHide Solution↓

(i) The complementary event S′ is that it will not snow tomorrow.

Hence, S′ is the event that it will not snow tomorrow.

(ii) We know that,

P(S) + P(S′) = 1

Given, P(S) = 0.03

P(S′) = 1 - P(S)

= 1 - 0.03

= 0.97

Hence, P(S′) = 0.97.

2Exercise 25(B)

Five students A, B, C, D and E are competing in a long distance race. Each student's probability of winning the race is given below:

A → 20%, B → 22%, C → 7%, D → 15% and E → 36%

(i) Who is most likely to win the race?

(ii) Who is least likely to win the race?

(iii) Find the sum of probabilities given.

(iv) Find the probability that either A or D will win the race.

(v) Let S be the event that B will win the race.

(a) Find P(S)

(b) State, in words, the complementary event S′

(c) Find P(S′)

Show SolutionHide Solution↓

The given probabilities are:

A → 20%, B → 22%, C → 7%, D → 15% and E → 36%.

(i) Since E has the greatest probability, E is most likely to win the race.

Hence, E is most likely to win the race.

(ii) Since C has the least probability, C is least likely to win the race.

Hence, C is least likely to win the race.

(iii) Sum of the given probabilities = 20% + 22% + 7% + 15% + 36%

= 100%

= 1

Hence, the sum of the given probabilities is 100%, i.e., 1.

(iv) Since only one student can win the race, the events that A wins and D wins are mutually exclusive.

P(A or D wins) = 20% + 15% = 35%

Hence, the probability that either A or D will win the race is 35%, i.e., 0.35.

(v) Let S be the event that B will win the race.

(a) P(S) = 22% = 0.22

Hence, P(S) = 0.22.

(b) The complementary event S′ is that B will not win the race.

Hence, S′ is the event that B will not win the race.

(c) We know that,

P(S) + P(S′) = 1

P(S′) = 1 - P(S)

= 1 - 0.22

= 0.78

Hence, P(S′) = 0.78, i.e., 78%.

3Exercise 25(B)

A ticket is randomly selected from a basket containing 3 green, 4 yellow and 5 blue tickets. Determine the probability of getting:

(a) a green ticket

(b) a green or yellow ticket

(c) an orange ticket

Show SolutionHide Solution↓

Total number of tickets = 3 + 4 + 5 = 12

(a) Number of favourable outcomes (Getting a green ticket) = 3

$$\begin{aligned}\text{P(Getting a green ticket)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{3}{12} \\[1em] = \frac{1}{4}\end{aligned}$$

Hence, the probability of getting a green ticket is $\mathbf{\frac{1}{4}}$.

(b) Number of green or yellow tickets = 3 + 4 = 7

Number of favourable outcomes (Getting a green or yellow ticket) = 7

$$\begin{aligned}\text{P(Getting a green or yellow ticket)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{7}{12}\end{aligned}$$

Hence, the probability of getting a green or yellow ticket is $\mathbf{\frac{7}{12}}$.

(c) There is no orange ticket in the basket.

Number of favourable outcomes (Getting an orange ticket) = 0

$$\begin{aligned}\text{P(Getting an orange ticket)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{0}{12} = 0\end{aligned}$$

Hence, the probability of getting an orange ticket is 0.

4Exercise 25(B)

Ten cards with numbers 1 to 10 written on them are placed in a bag. A card is chosen from the bag at random. Determine the probability of choosing:

(a) 7

(b) 9 or 10

(c) a number greater than 4

(d) a number less than 6

Show SolutionHide Solution↓

Total number of cards = 10

(a) Favourable outcomes for choosing 7 = 7

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Choosing 7)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{10}\end{aligned}$$

Hence, the probability of choosing 7 is $\mathbf{\frac{1}{10}}$.

(b) Favourable outcomes for choosing 9 or 10 = 9, 10

No. of favourable outcomes = 2

$$\begin{aligned}\text{P(Choosing 9 or 10)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{2}{10} \\[1em] = \frac{1}{5}\end{aligned}$$

Hence, the probability of choosing 9 or 10 is $\mathbf{\frac{1}{5}}$.

(c) The numbers greater than 4 are 5, 6, 7, 8, 9 and 10.

Number of favourable outcomes (Choosing a number greater than 4) = 6

$$\begin{aligned}\text{P(Choosing a number greater than 4)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{6}{10} \\[1em] = \frac{3}{5}\end{aligned}$$

Hence, the probability of choosing a number greater than 4 is $\mathbf{\frac{3}{5}}$.

(d) The numbers less than 6 are 1, 2, 3, 4 and 5.

Number of favourable outcomes (Choosing a number less than 6) = 5

$$\begin{aligned}\text{P(Choosing a number less than 6)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{5}{10} \\[1em] = \frac{1}{2}\end{aligned}$$

Hence, the probability of choosing a number less than 6 is $\mathbf{\frac{1}{2}}$.

5Exercise 25(B)

A carton contains eight brown and four white eggs. Find the probability that an egg selected at random is:

(a) brown

(b) white

Show SolutionHide Solution↓

Total number of eggs = 8 + 4 = 12

(a) Number of favourable outcomes (Selecting a brown egg) = 8

$$\begin{aligned}\text{P(Selecting a brown egg)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{8}{12} \\[1em] = \frac{2}{3}\end{aligned}$$

Hence, the probability of selecting a brown egg is $\mathbf{\frac{2}{3}}$.

(b) Number of favourable outcomes (Selecting a white egg) = 4

$$\begin{aligned}\text{P(Selecting a white egg)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{4}{12} \\[1em] = \frac{1}{3}\end{aligned}$$

Hence, the probability of selecting a white egg is $\mathbf{\frac{1}{3}}$.

6Exercise 25(B)

A box contains 3 yellow, 4 green and 8 blue tickets. A ticket is chosen at random. Find the probability that the ticket is:

(a) yellow

(b) green

(c) blue

(d) red

(e) not yellow

Show SolutionHide Solution↓

Total number of tickets = 3 + 4 + 8 = 15

(a) Number of favourable outcomes (Getting a yellow ticket) = 3

$$\begin{aligned}\text{P(Getting a yellow ticket)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{3}{15} \\[1em] = \frac{1}{5}\end{aligned}$$

Hence, the probability of getting a yellow ticket is $\mathbf{\frac{1}{5}}$.

(b) Number of favourable outcomes (Getting a green ticket) = 4

$$\begin{aligned}\text{P(Getting a green ticket)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{4}{15}\end{aligned}$$

Hence, the probability of getting a green ticket is $\mathbf{\frac{4}{15}}$.

(c) Number of favourable outcomes (Getting a blue ticket) = 8

$$\begin{aligned}\text{P(Getting a blue ticket)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{8}{15}\end{aligned}$$

Hence, the probability of getting a blue ticket is $\mathbf{\frac{8}{15}}$.

(d) There is no red ticket in the box.

Number of favourable outcomes (Getting a red ticket) = 0

$$\begin{aligned}\text{P(Getting a red ticket)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{0}{15} = 0\end{aligned}$$

Hence, the probability of getting a red ticket is 0.

(e) Number of tickets which are not yellow = 4 + 8 = 12

Number of favourable outcomes (Getting a ticket which is not yellow) = 12

$$\begin{aligned}\text{P(Getting a ticket which is not yellow)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{12}{15} \\[1em] = \frac{4}{5}\end{aligned}$$

Hence, the probability of getting a ticket which is not yellow is $\mathbf{\frac{4}{5}}$.

7Exercise 25(B)

The following table shows number of males and number of females of a small locality in different age groups.

Age in yearsMaleFemale
10-2086
21-501210
Above 5064

If one of the persons, from this locality, is picked at random, what is the probability that

(a) the person picked is a male?

(b) the person picked is a female?

(c) the person picked is a female aged 21-50?

(d) the person picked is a male with age upto 50 years?

Show SolutionHide Solution↓

Number of males = 8 + 12 + 6 = 26

Number of females = 6 + 10 + 4 = 20

Total number of persons = 26 + 20 = 46

(a) Number of favourable outcomes (Picking a male) = 26

$$\begin{aligned}\text{P(Picking a male)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{26}{46} \\[1em] = \frac{13}{23}\end{aligned}$$

Hence, the probability that the person picked is a male is $\mathbf{\frac{13}{23}}$.

(b) Number of favourable outcomes (Picking a female) = 20

$$\begin{aligned}\text{P(Picking a female)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{20}{46} \\[1em] = \frac{10}{23}\end{aligned}$$

Hence, the probability that the person picked is a female is $\mathbf{\frac{10}{23}}$.

(c) Number of females aged 21–50 years = 10

$$\begin{aligned}\text{P(Picking a female aged 21–50 years)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{10}{46} \\[1em] = \frac{5}{23}\end{aligned}$$

Hence, the probability that the person picked is a female aged 21–50 years is $\mathbf{\frac{5}{23}}$.

(d) Number of males aged up to 50 years = 8 + 12 = 20

$$\begin{aligned}\text{P(Picking a male aged up to 50 years)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{20}{46} \\[1em] = \frac{10}{23}\end{aligned}$$

Hence, the probability that the person picked is a male aged up to 50 years is $\mathbf{\frac{10}{23}}$.

8Exercise 25(B)

A coin is tossed 400 times and we get head 270 times. If the same coin is tossed one more time, find the probability that this time the coin comes out to be:

(a) head

(b) tail

Show SolutionHide Solution↓

Total number of trials = 400

Number of times head is obtained = 270

Number of times tail is obtained = 400 − 270 = 130

(a) Number of favourable outcomes (Getting a head) = 270

$$\begin{aligned}\text{P(Getting a head)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{270}{400} \\[1em] = \frac{27}{40}\end{aligned}$$

Hence, the probability of getting a head is $\mathbf{\frac{27}{40}}$.

(b) Number of favourable outcomes (Getting a tail) = 130

$$\begin{aligned}\text{P(Getting a tail)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{130}{400} \\[1em] = \frac{13}{40}\end{aligned}$$

Hence, the probability of getting a tail is $\mathbf{\frac{13}{40}}$.

9Exercise 25(B)

A die is rolled 8 times and the number 5 is obtained 2 times. Now, if the same die is rolled again, find the probability of getting a 5.

Show SolutionHide Solution↓

Total number of trials = 8

Number of times 5 is obtained = 2

$$\begin{aligned}\text{P(Getting 5)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{2}{8} \\[1em] = \frac{1}{4}\end{aligned}$$

Hence, the probability of getting 5 is $\mathbf{\frac{1}{4}}$.

10Exercise 25(B)

A die is rolled once. Find the probability of getting:

(a) 3

(b) 5

(c) 2 or 4

(d) more than 4

Show SolutionHide Solution↓

In a single roll of a die, the possible outcomes are 1, 2, 3, 4, 5 and 6.

Total number of possible outcomes = 6

(a) Favourable outcomes for getting 3 = 3

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Getting 3)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{6}\end{aligned}$$

Hence, the probability of getting 3 is $\mathbf{\frac{1}{6}}$.

(b) Favourable outcomes for getting 5 = 5

No. of favourable outcomes = 1

$$\begin{aligned}\text{P(Getting 5)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{1}{6}\end{aligned}$$

Hence, the probability of getting 5 is $\mathbf{\frac{1}{6}}$.

(c) Favourable outcomes for getting 2 or 4 = 2, 4

No. of favourable outcomes = 2

$$\begin{aligned}\text{P(Getting 2 or 4)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{2}{6} \\[1em] = \frac{1}{3}\end{aligned}$$

Hence, the probability of getting 2 or 4 is $\mathbf{\frac{1}{3}}$.

(d) The numbers more than 4 are 5 and 6.

Number of favourable outcomes (Getting a number more than 4) = 2

$$\begin{aligned}\text{P(Getting a number more than 4)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{2}{6} \\[1em] = \frac{1}{3}\end{aligned}$$

Hence, the probability of getting a number more than 4 is $\mathbf{\frac{1}{3}}$.

11 Mark · Multiple Choice Questions

A die is tossed once, the probability of getting a number greater than 4 is:

Show SolutionHide Solution↓

In a single throw of a die, the possible outcomes are 1, 2, 3, 4, 5 and 6.

Total number of possible outcomes = 6

The numbers greater than 4 are 5 and 6.

Number of favourable outcomes = 2

$$\begin{aligned}\text{P(Getting a number greater than 4)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{2}{6} \\[1em] = \frac{1}{3}\end{aligned}$$

Hence, option 3 is the correct option.

21 Mark · Multiple Choice Questions

A die is tossed once, the probability of getting a number greater than 6 is:

Show SolutionHide Solution↓

In a single throw of a die, the possible outcomes are 1, 2, 3, 4, 5 and 6.

There is no number greater than 6 on a die.

Number of favourable outcomes = 0

$$\begin{aligned}\text{P(Getting a number greater than 6)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{0}{6} = 0\end{aligned}$$

Hence, option 4 is the correct option.

31 Mark · Multiple Choice Questions

A bag contains 5 red balls, 8 blue balls and 7 green balls. All balls are identical in size. One ball is drawn at random from the bag, the probability of getting a blue ball is:

Show SolutionHide Solution↓

Total number of balls = 5 + 8 + 7 = 20

Number of favourable outcomes (Getting a blue ball) = 8

$$\begin{aligned}\text{P(Getting a blue ball)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{8}{20} \\[1em] = \frac{2}{5}\end{aligned}$$

Hence, option 1 is the correct option.

41 Mark · Multiple Choice Questions

The probability of an impossible event is:

Show SolutionHide Solution↓

An impossible event has no favourable outcome.

No. of favourable outcomes = 0

$$\begin{aligned}\text{P(Impossible event)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{0}{\text{Total number of possible outcomes}} = 0\end{aligned}$$

Therefore, the probability of an impossible event is 0.

Hence, option 2 is the correct option.

51 Mark · Multiple Choice Questions

Probability of selecting a vowel from the word MUSSOORIE is:

Show SolutionHide Solution↓

The word MUSSOORIE has 9 letters: M, U, S, S, O, O, R, I and E.

Total number of possible outcomes = 9

The vowels are U, O, O, I and E.

Number of favourable outcomes (Selecting a vowel) = 5

$$\begin{aligned}\text{P(Selecting a vowel)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{5}{9}\end{aligned}$$

Hence, option 4 is the correct option.

61 Mark · Multiple Choice Questions

A die is tossed once, the probability of getting 3 and 5 is:

Show SolutionHide Solution↓

In one throw of a die, only one number can appear.

Therefore, getting both 3 and 5 in the same throw is impossible.

Total number of possible outcomes = 6

No. of favourable outcomes = 0

$$\begin{aligned}\text{P(Getting 3 and 5)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{0}{6} = 0\end{aligned}$$

Hence, option 2 is the correct option.

71 Mark · Multiple Choice Questions

A die is tossed once, probability of getting 3 or 5 is:

Show SolutionHide Solution↓

In a single throw of a die, the possible outcomes are 1, 2, 3, 4, 5 and 6.

Total number of possible outcomes = 6

Favourable outcomes for getting 3 or 5 = 3, 5

No. of favourable outcomes = 2

$$\begin{aligned}\text{P(Getting 3 or 5)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{2}{6} \\[1em] = \frac{1}{3}\end{aligned}$$

Hence, option 1 is the correct option.

81 Mark · Multiple Choice Questions

A card is drawn from a pack of 50 identical cards, numbered from 1 to 50. The probability of drawing a square number is:

Show SolutionHide Solution↓

The square numbers from 1 to 50 are 1, 4, 9, 16, 25, 36 and 49.

Total number of possible outcomes = 50

Number of favourable outcomes (Drawing a square number) = 7

$$\begin{aligned}\text{P(Drawing a square number)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{7}{50}\end{aligned}$$

Hence, option 3 is the correct option.

91 Mark · Multiple Choice Questions

There are 12 thin cards, all of the same size. Names of 12 months of the year are written on these cards (one month on one card). If one card is drawn at random, the probability of getting the name of the month starting from J is:

Show SolutionHide Solution↓

The months whose names start with J are January, June and July.

Total number of possible outcomes = 12

Number of favourable outcomes = 3

$$\begin{aligned}\text{P(Getting the name of a month starting with J)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{3}{12} \\[1em] = \frac{1}{4}\end{aligned}$$

Hence, option 1 is the correct option.

101 Mark · Multiple Choice Questions

A die is thrown 100 times and the outcomes are noted as shown below:

OutcomesNo. of times
120
215
315
420
520
610

If the same die is thrown again, the probability of getting a number less than 4 is:

Show SolutionHide Solution↓

Total number of outcomes = 100

The numbers less than 4 are 1, 2 and 3.

Number of times 1, 2 or 3 appears = 20 + 15 + 15 = 50

Number of favourable outcomes = 50

$$\begin{aligned}\text{P(Getting a number less than 4)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{50}{100} \\[1em] = \frac{1}{2}\end{aligned}$$

Hence, option 2 is the correct option.

11Statement I-II Type Questions

Statement 1: A die is thrown once. The probability of getting a prime number greater than six is 0.

Statement 2: If there is no chance that an event may occur, such events are called impossible events.

Which of the following options is correct?

Show SolutionHide Solution↓

For Statement 1:

In a single throw of a die, the possible outcomes are 1, 2, 3, 4, 5 and 6.

There is no prime number greater than 6 among these outcomes.

Total number of possible outcomes = 6

No. of favourable outcomes = 0

$$\begin{aligned}\text{P(Getting a prime number greater than 6)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{0}{6} = 0\end{aligned}$$

Therefore, getting a prime number greater than 6 is an impossible event and its probability is 0.

So, Statement 1 is true.

For Statement 2:

An event which has no chance of occurring is called an impossible event.

So, Statement 2 is true.

∴ Both the statements are true.

Hence, option 1 is the correct option.

12Statement I-II Type Questions

Statement 1: A random experiment is that in which all possible outcomes are not known.

Statement 2: Probability of an event can be expressed as a percentage between 0% to 100%.

Which of the following options is correct?

Show SolutionHide Solution↓

For Statement 1:

In a random experiment, the possible outcomes are known, but it is not known in advance which one of them will occur.

Therefore, Statement 1 is false.

For Statement 2:

The probability of an event lies between 0 and 1, both inclusive. Hence, it can also be expressed as a percentage from 0% to 100%.

Therefore, Statement 2 is true.

∴ Statement 1 is false, and Statement 2 is true.

Hence, option 4 is the correct option.

13Assertion-Reason Type Questions

Assertion (A): The probability of getting an odd number from first 10 even numbers = $\frac{1}{2}$.

Reason (R): Let an event is represented by P(A), then P(A) = $\frac{\text{Number of favourable out comes}}{\text{Total number of possible outcomes}}$

Show SolutionHide Solution↓

The first 10 even numbers are 2, 4, 6, 8, 10, 12, 14, 16, 18 and 20.

There is no odd number among them.

Therefore,

$$\begin{aligned}\text{P(Getting an odd number)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}} \\[1em] = \frac{0}{10} = 0\end{aligned}$$

So, Assertion (A) is false.

The probability of an event A is given by

$\text{P(A)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}}$

So, Reason (R) is true.

∴ A is false and R is true.

Hence, option 2 is the correct option.

14Assertion-Reason Type Questions

Assertion (A): Three students P, Q and R attempting to solve a mathematics problem. Their respective probabilities of solving is given as: P → 32%, Q → 36%, R → 32%.

When asked by the teacher, other students respond that Q is most likely to solve that problem.

Reason (R): Out of all given probabilities, the events with equal probabilities are said to be most likely events.

Show SolutionHide Solution↓

The probabilities of solving the problem are:

P → 32%, Q → 36% and R → 32%.

Since 36% is the greatest probability, Q is most likely to solve the problem.

So, Assertion (A) is true.

Events having equal probabilities are called equally likely events. They are not necessarily the most likely events.

So, Reason (R) is false.

∴ A is true and R is false.

Hence, option 1 is the correct option.

15Assertion-Reason Type Questions

Assertion (A): The chance of a newly born baby to be girl is 0.5.

Reason (R): If the chances of happening and non-happening of an event are same, we say that event has even chances of happening.

Show SolutionHide Solution↓

Assuming that the chances of a newly born baby being a girl or a boy are equal,

$$\begin{aligned}\text{P(A baby is a girl)} = \frac{\text{Number of favourable outcomes}}{\text{Total number of outcomes}} \\[1em] = \frac{1}{2} = 0.5\end{aligned}$$

So, Assertion (A) is true.

If the chances of happening and non-happening of an event are the same, the event is said to have even chances of happening.

So, Reason (R) is true.

∴ Both A and R are true.

Hence, option 3 is the correct option.

Press ↓ to reveal a solution and move to the next question. Back to Selina Solutions Concise Mathematics Class 7 ICSE.